OpenStudy (anonymous):

Three variables x,y,z exist such that x^2+y^2+z^2 = 50 , xy + yz + zx = 47 ; xyz = 60 (i) (x+y+z)^3 = ? -- {i know this one!} (ii) What's the value of x ? (iii) What's the value of y ?

OpenStudy (anonymous):

This is a system of equations. First, you need to find the value of x, y, and z. I recommend doing this by using the last equation: xyz = 60 x = 60/yz Then you can substitute that in and find the value of y, z, and x.

OpenStudy (anonymous):

@thechocoluver445 : I did the very same thing.. but at the end, i got an equation, that i wasnt able to solve.. and that ruined my whole work! :/

OpenStudy (anonymous):

(x+y+z)^2= x^2+y^2+z^2+2(xy + yz + zx)=50+2*47=154 so we have x+y+z=sqrt(154) or x+y=sqrt(154)-z (I) xyz=60 so xy=60/z (II) xy + yz + zx=47 from (I) and (II) xy + yz + zx=xy+z(x+y)=60/z+z(sqrt(154)-z)=47 then u may solve \[-z^3+\sqrt{154} \ z^2-47z+60=0\]

OpenStudy (anonymous):

for solving such equations http://www.numberempire.com/equationsolver.php

OpenStudy (anonymous):

@mukushla : Dear, you probably did the steps correctly , but x+y+z = 12.. as in that first answer what you got as 154 , is actually 144 ..

OpenStudy (anonymous):

oh much better

OpenStudy (anonymous):

(x+y+z)^2= x^2+y^2+z^2+2(xy + yz + zx)=50+2*47=144 so we have x+y+z=12 or x+y=12-z (I) xyz=60 so xy=60/z (II) xy + yz + zx=47 from (I) and (II) xy + yz + zx=xy+z(x+y)=60/z+z(12-z)=47 we must solve \[−z^3+12z^2−47z+60=0\]

OpenStudy (anonymous):

thats very interesting the solutions for z are z=3 z=4 z=5

OpenStudy (anonymous):

(x,y,z)=(3,4,5)

OpenStudy (anonymous):

Exactly .. !! Thanx to that "equation solver" of yours.. Helped me a ton! :))